### v2

# Neither Stochastic Parroting nor AGI: LLMs Solve Tasks through Context-Directed Extrapolation from Training Data Priors

Abstract
In this position paper we raise critical awareness of a realistic view of LLM capabilities that eschews extreme alternative views that LLMs are either ‘stochastic parrots’ or in possession of ‘emergent’ advanced reasoning capabilities, which, due to their unpredictable emergence, constitute an existential threat. Our middle-ground view is that LLMs extrapolate from priors from their training data while using context to guide the model to the appropriate priors; we call this ‘context-directed extrapolation.’ Specifically, this context direction is achieved through examples in base models, leading to in-context learning, while instruction tuning allows LLMs to perform similarly based on prompts rather than explicit examples. Under this view, substantiated though existing literature, while reasoning capabilities go well beyond stochastic parroting, such capabilities are predictable, controllable, not indicative of advanced reasoning akin to high-level cognitive capabilities in humans, and not infinitely scalable with additional training. As a result, fears of uncontrollable emergence of agency are allayed, while research advances are appropriately refocused on the processes of context-directed extrapolation and how this interacts with training data to produce valuable capabilities in LLMs. Future work can therefore explore alternative augmenting techniques that do not rely on inherent advanced reasoning in LLMs.
1 Introduction
Large Language Models (LLMs), primarily trained on the cloze-inspired task of next-token prediction, have demonstrated the ability to solve a range of tasks, some of which typically require reasoning in humans [11, 62, 52]. Understanding the basis for how LLMs achieve this is central to several key considerations, including the continued development of these models, their safety and security implications, and the contexts in which they should or should not be used [25]. Simultaneously, better understanding the underlying framework used by LLMs to solve such tasks has significant implications for various research domains. For example, it would suggest domains wherein the use of LLMs would require additional guardrails and training.Therefore, providing a clear explanation for LLM performance is crucial. However, despite, or perhaps because of, its importance, there has been little consensus on how LLMs are able to solve complex tasks. Currently, two predominant viewpoints exist: the first (alternative view 1) is that LLMs are essentially ‘stochastic parrots,’ predicting statistically likely outputs in line with their training objectives [6, 5, 37]. The second viewpoint (alternative view 2) posits that as LLMs are scaled up, in terms of both parameters and training data, they develop ‘emergent reasoning’ capabilities [11, 62, 52], which, in some cases, have been described as “sparks of AGI” (Artificial General Intelligence) [13]. It is important to note that our discussion focuses on capabilities that are either already at the level of, or indicative of the potential for, advanced general-purpose reasoning akin to high level human cognitive capabilities. For brevity we will call this ‘advanced reasoning.’
Neither of these viewpoints, we will argue, sufficiently explain the capabilities and limitations of state-of-the-art LLMs. Specifically, LLMs trained only on next token prediction can solve a range of non-memorizable tasks that require more than generating the next most likely token (see Section 3). Nonetheless, these models fail on a range of tasks which humans are able to trivially solve (Section 4), indicating that they are not capable of certain types of human-like reasoning. Note that our discussion of reasoning focuses on capabilities that are either already at the level of, or indicative of the potential for, advanced general-purpose reasoning akin to high level human cognitive capabilities.
Despite the conflicting findings and viewpoints surrounding the explanation of LLM performance, there is one ability that LLMs possess indisputably: this is the ability of LLMs, trained solely on next-token prediction (referred to as Base LLMs), to solve a range of tasks when provided with a few examples in the input prompt—an ability known as in-context learning (ICL) or few-shot learning [12, 40]. For example, an LLM provided with the input prompt “1 + 3 = 5; 7 + 12 = 20; 8 + 3 =” will respond with ‘12.’ Notice that this requires the model to identify this task to be a form of modified addition ($a+b+1$) and perform the required modified calculation. [^1] We test this on LLaMA 3 70B Base
The idea that LLMs are just “stochastic parrots” is challenged by their ability to learn in-context. For example, research by Bonial et al. [10] shows that LLMs can be taught entirely novel phrases, or Multi-Word Expressions (MWEs), on the fly. After seeing just one example defining “winking at pringles,” a novel MWE assigned the meaning ‘to indulge in frivolity,’ models could correctly interpret and even reason about new uses of the phrase. Since these novel MWEs could not have been in the training data, this performance demonstrates a capability beyond simply predicting the next most likely word.
However, the view that LLMs are stochastic parrots is not without some merit. LLMs, despite their self-supervised training objectives, are fundamentally machine learning systems. As such, it is reasonable to assume that they operate in the same way as other machine learning models: by solving tasks using priors extrapolated from their training data. Indeed, after the pre-training process, LLMs are fine-tuned with instructions [61] to enable them to follow directives and align with human preferences [42]. Thus, the responses LLMs generate result from the combination of priors extrapolated from all aspects of their training. Therefore, LLMs, like all machine learning systems, do generate output by relying on priors extracted from their training data, whether from pre-training, instruction tuning, or safety alignment.
There is, however, one important caveat: unlike traditional machine learning systems, LLMs have in-context capabilities. Therefore, we propose that extrapolation and in-context learning together should serve as the criteria for explaining the performance of LLMs, unless evidence to the contrary emerges, which, as we will demonstrate, has not yet been presented. Importantly, since instruction-tuned LLMs can respond to prompts without explicit examples, we propose that the instruction-tuning process enables models to target relevant priors without the need for such examples. Indeed, some works call this ability of instruction-tuned models to leverage contextual information for task-solving ‘in-context learning’ [15].
We will argue that context-directed extrapolation from training data, exemplified by ICL in Base models, can give the impression of stochastic parroting when the target task closely resembles the training data, and that it can also give the impression of advanced reasoning when the target task is clearly extrapolable from training data. Context guided extrapolation from pre-training data can also explain other peculiar characteristics of LLMs, such as their tendency to generate fluent text that is not in line with reality, known as hallucinations. Specifically, when the extrapolation from pre-training data leads to an answer outside of the correct range, hallucinations are produced.
Our position: As opposed to stochastic parroting, or advanced general-purpose reasoning akin to high-level human cognitive capabilities, we argue that context-directed extrapolation from training data priors offers a comprehensive framework to adequately explain the capabilities and limitations of LLMs. This context direction is achieved through examples in base models, leading to in-context learning, while instruction tuning allows LLMs to perform similarly based on prompts rather than explicit examples.
Significance: The framework used to interpret the capabilities of LLMs directly influences the impact of this rapidly evolving field. Dismissing LLMs as stochastic parrots is just as counterproductive as attributing them with advanced reasoning akin to human-level cognitive capabilities. Specifically, dismissing their capabilities could slow the adoption of these models in applications where their use can enable more flexible natural-language interaction with computational systems. Equally, claiming that these models exhibit advanced reasoning or AGI-like abilities could undermine public trust. This rhetoric has already fueled discussions about the existential threats [2] posed by LLMs and their potential for deception [22]. Such discussions could stifle research in this area and, if human-level general reasoning is indeed the goal for some researchers, could hinder progress in that direction. Instead of relegating this discussion to the fringes, we believe that bringing it to the forefront of academic discourse is essential, especially at a time when LLM research is rapidly expanding. Raising awareness about our realistic, middle-ground view that LLMs reasoning capabilities are predictable and controllable appropriately refocuses safety and research questions on understanding all we can about context-directed extrapolation while reducing fears of uncontrollable emergence of agency. Conversely, failing to recognize this intermediate viewpoint could significantly and negatively impact the future direction of research, public trust, and policymaking.
2 Defining Reasoning
We embrace a working definition of reasoning based on the cognitive processes outlined in Krathwohl [31], who revises Bloom’s original taxonomy of educational objectives [9]. Although the taxonomy’s original purpose was to provide a framework for testing educational objectives, it is also a well-recognized rubric for defining and categorizing levels of mastery within the cognitive domain generally, and has been leveraged within AI as well (e.g., Blaha et al. [8]). Like the original taxonomy, Krathwohl’s revised taxonomy breaks up the categories across the dimensions of Cognitive Process and Knowledge.
In this research, we focus on the Cognitive Process dimension, which is an ordered set (from simple to more complex) of processes that subdivide reasoning: Remember, Understand, Apply, Analyze, Evaluate, Create. Within these processes, we focus on Remember, Understand and Apply. Our focus is guided by the fact that this is considered an ordered set of processes, such that the higher-level processes of Evaluate and Create rely upon mastery of the lower-level cognitive processes; within this classification we find evidence that LLMs are limited to Understand, which would preclude higher-order processes. [^2] We also caution the reader against assuming commonplace definitions of these terms—there are certainly levels of “understanding” that models do not achieve, and it is clear that models can do some types of evaluation and can create new texts, but this does not necessarily meet the definitional criteria of these terms as used in educational settings. Remember is defined as “Retrieving relevant knowledge from long-term memory,” and is split into subcategories of Recognizing and Recalling [31, p. 215]. Understand is defined as “Determining the meaning of instructional messages, including oral, written, and graphic communication,” and is split into finer grained categories of Interpreting, Exemplifying, Classifying, Summarizing, Inferring, Comparing, and Explaining (ibid.). Apply is defined as “Carrying out or using a procedure in a given situation,” but we also leverage Bloom’s original Application description, which involves solving a new, unseen problem by deploying and operationalizing previously learned information. Apply is split into Executing and Implementing.
Because we argue that LLM reasoning is limited to Understand and does not reach the level of Apply, we will further explain the distinction between these two processes. If we assume a set of items that have been labeled as being members of a certain class, then Understanding this information assumes that one can explain why the items are members of that class, as well as compare the known set items to other items in order to further classify items and provide examples of the class. This is similar to teaching a student about the parts of speech of their language, then asking the student to explain what a verb is, provide some examples of verbs, and recognize whether or not a familiar word is a verb. To engage in the Apply process, one must first Understand, but then also determine how to deploy that understanding with respect to a new set of items where the generalization that makes this set a class is not already given. This is similar to asking a linguist to examine the morpho-syntax of a new language and determine what the parts of speech are in that language—the linguist will generalize knowledge of the features shared by members of part-of-speech classes in other languages and apply this knowledge to the new language.
In terms of this theoretical framework, we argue that LLM reliance on pretraining data enables -Remembering of Factual Knowledge, which necessarily requires memorization as opposed to a deeper level of the Cognitive Process dimension. Furthermore, LLMs certainly Remember an immense amount of Conceptual Knowledge, including the memorization of theories and even potentially commonplace generalizations; we emphasize that such Remembering may, to the human eye, masquerade as Understanding and even Applying, as people are more likely to apply a heuristic in lieu of memorizing web-scale information. Thus, we may attribute similar cognitive processes to LLMs where they are in fact able to rely on memory.
Nonetheless, in effectively extrapolating [^3] We note that the term extrapolation is also sometimes applied to human cognitive processes; however, as described in that literature, it is a more advanced reasoning process than mathematical extrapolation. For example, Aviles [3] describes extrapolation as a part of Understanding (termed ‘Comprehension’ in Bloom’s original taxonomy), and indicates that it involves identifying trends and consequences (e.g., describing the intended and unintended consequences of social policies). Within AI, identifying consequences is considered more advanced causal reasoning, and, in the case of a potential social policy, this would be akin to counterfactual reasoning, as opposed to extrapolation [43]. from immense amounts of pretraining data, LLMs are able to demonstrate Understanding of knowledge that is well-represented in pretraining data. Extrapolation from pretraining data is a combination of Remembering, Understanding and hallucination. We argue that when the pretraining data closely resembles the target task, memory is sufficient for what is essentially stochastic parroting. When the prompt includes contextual direction, either through ICL examples in base models, or through the prompt itself in instruction tuned models, then models are able to effectively extrapolate from their pretraining data, demonstrating some Understanding as defined above. Finally, if the pretraining data to be extrapolated from is not effectively triggered by the prompt (either examples or directly through the prompt) or that data simply is not present, models will hallucinate. While not factually correct, hallucinations may still maintain the illusion of deeper cognitive processes, especially to the human eye, which is likely to attribute human-like reasoning (the ‘ELIZA effect’ [66]).
Overall, we argue and provide evidence that LLM reasoning is limited to what is termed Understanding in Bloom’s framework. We also argue that we cannot test whether or not the reasoning process of Apply takes place unless i) we can guarantee that Remembering is not the likely explanation given the relevance of pretraining data to answering examples in the test set, and ii) that successful performance on evaluations cannot be attributed to contextual-guided extrapolation, which simplifies the task to the basic Understanding processes, including Classifying (the inverse of Exemplifying), as well as Comparing test items to the provided examples.
3 Alternative View: Stochastic Parroting
The notion that LLMs are ‘Stochastic Parrots’ [^4] We note that stochastic parroting does not have a precise parallel in the literature on human cognitive processes described in Section 2. It relies strongly on Remembering, but rather than memory of chunks of information, it also involves extrapolation to guess the next most likely token, based upon what is in memory. Here, the term extrapolation refers to the mathematical, machine-learning sense: estimating values outside the range of known data points based on the observed relationship between variables. that respond with the next most likely token was first introduced by Bender et al. [6]. They rightly argue that statistical generalization from web data at scale can lead to the propagation of biases inherent to the statistical patterns present in the text on the web. We do not contradict this argument. However, Bender et al. [6] also suggests that LLMs merely generate the next most likely token.
As shown by the novel MWE example in Section 1, this claim is demonstrably false. In the novel MWE interpretation task of Bonial et al. [10], a stochastic parrot generating statistically probable outputs would have no existing knowledge of the assigned phrasal meaning of “winking at pringles.” Rather, it would only have access to information about the meanings of the individual lexical items as used in pretraining data. It would therefore likely produce a literal or nonsensical completion. The fact that models can use the definition provided in the prompt to ‘learn’ the MWE’s novel meaning (to indulge in frivolity) and correctly reason about it in new situations provides clear evidence against the notion that they are just stochastic parrots.
Nonetheless, the assumption that LLMs do nothing but generate the next most likely token itself should not be dismissed out of hand. After all, this is what they are trained to do [46], and it is reasonable to expect them to behave in this manner. Indeed, LLMs do exhibit hallucinations [26], and this significant shortcoming does suggest that they might only be generating the next most likely token.
At the same time, LLMs are trained on vast amounts of data, much of which is often confidential, making it difficult to assess them on truly novel problems that they have not seen before [47]. Additionally, there is ample evidence suggesting that LLMs sometimes take shortcuts in solving tasks [71]. There have also been instances of data leakage [75], where evaluation data may have inadvertently been included in the pre-training or fine-tuning datasets. Furthermore, some datasets used to evaluate LLMs may not adequately test reasoning abilities [36], instead allowing models to solve them based on model’s access to vast memory. While humans must reason to solve some of these tasks due to limited memory, LLMs may not need to reason in the same way, as they have access to far more extensive memory. These criticisms are valid, but they do not imply that all datasets are tainted by such issues.
Furthermore, it is now undeniable that sufficiently large LLMs trained on sufficiently large corpora on the next-token prediction task exhibit in-context capabilities or the ability to solve tasks based on a few examples, as described above [12, 62, 53, 36, 65].
For example, while LLMs trained solely on next-token prediction (i.e. base models) cannot solve tasks that require abstract reasoning, they can solve such problems when provided with examples through ICL [36]. An illustration of this is the following logical deduction problem from the Big-Bench benchmark:
Question: On a shelf, there are five books: a gray book, a red book, a purple book, a blue book, and a black book. The red book is to the right of the gray book. The black book is to the left of the blue book. The blue book is to the left of the gray book. The purple book is the second from the right.
Targets: ‘The gray book is the leftmost.’: 0; ‘The red book is the leftmost.’: 0; ‘The purple book is the leftmost.’: 0; ‘The blue book is the leftmost.’: 0; ‘The black book is the leftmost.’: 1
The difference in performance when base models are provided with ICL examples and when they are not offers clear evidence for our framework of context-directed extrapolation. Base models fail on such reasoning tasks because they require explicit examples in the prompt to provide the ‘context’ needed to guide their extrapolation. In contrast, instruction-tuned models can derive this context directly from the prompt, allowing them to solve the problem without needing any examples, as suggested by Lu et al. [36].
Finally, the argument against ‘stochastic parroting’ becomes even stronger when you use novel words, which forces a model to reason instead of just relying on statistical patterns. In addition to the novel MWE experiments presented above, the Winodict benchmark [20] provides a similar test case by adapting the Winograd Schema Challenge [32], which offers a classic method for evaluating a system’s capacity for common-sense reasoning through the task of anaphora resolution. The original Winograd task consists of sentence pairs that differ by only a single word, a change that completely alters the referent of an ambiguous pronoun. For example, consider ‘The man could not lift his son because he was so weak.’ and ‘The man could not lift his son because he was so heavy.’ In the first sentence, ‘he’ refers to the man and in the second, ‘he’ refers to the son.
The Winodict benchmark replaces the critical content word with a "nonce" word whose definition is provided entirely within the prompt. This transforms the task from a test of stored world knowledge to one of dynamic, in-context semantic application. For instance, a model is presented with the following: The verb ‘to plest’ means to be scared of. The city councilmen refused the demonstrators a permit because they plested violence. Here, the resolution of the pronoun ‘they’ is entirely dependent on the model’s ability to parse the ad-hoc definition of ‘plest’ and apply it to the sentence’s causal structure.
These dynamic, in-context capabilities truly set LLMs apart and make LLMs fundamentally different from all previous machine learning systems and challenge the notion of them being stochastic parrots.
4 Alternative View: Advanced Reasoning
In contrast to the notion of stochastic parroting, there has been a body of research showing that with increased scale, there is a clear increase in model abilities that can be argued to be indicative of capabilities akin to high-level cognitive reasoning processes in humans [12, 62, 53, 36, 65].
As we will demonstrate below, evidence of the kind of high-level cognitive reasoning that we are interested in stems only from models’ ability to solve tasks without being explicitly trained for them (this is what separates the more basic Understanding cognitive process from higher-level processes). While we will explore this in more detail in Section 5.2.1, it is important to note here that if a model has been specifically trained to perform a task, such as playing chess, and can outperform humans, it is not considered to possess the type of general, advanced reasoning capabilities we are concerned with as it does not involve the generalization of knowledge acquired with respect to one set of items to an entirely new set of items [16].
Two key elements in prior research discourse have been cited as evidence of advanced reasoning in LLMs. The first is the scaling laws, which show that larger models trained on more data are capable of solving a greater number of problems [28]. The second is the claim that sufficiently large LLMs exhibit what are known as emergent abilities [62]. We will address how both of these phenomena can be explained through our framework of extrapolation from pre-training data guided by ICL in Section 5. In this section, however, we will focus on evaluation-based evidence that LLMs are not always capable of performing high-level reasoning.
4.1 Failure on Relatively Simple Tasks
The first indication that LLMs are not performing advanced reasoning comes from direct tests of their performance. In fact, it has been shown that LLMs fail on a range of tasks that are relatively simple, especially when compared to the advanced capabilities they are often claimed to possess [38]. We will focus on one such task, which is so basic that children can solve it. Shapira et al. [51] test a variety of models on an established clinical psychology test administered to children that asks them to recognize and pinpoint a faux-pas made in a social situation, for example stating to a party host that one does not enjoy apple pie without realizing that this is precisely what the host has prepared for dessert. The authors find that the best-scoring models on this task (Flan-T5-xl and -xxl) obtain a score of 0.4 on the task, whereas children ages 9-11 score, on average, 0.82 (i.e. children can pinpoint the faux pas in twice as many short anecdotes as models).
There is also research attempting to use LLMs as AI planners for robots. Most of this research has been limited to the relatively simple Blocksworld domain, where planning involves taking in a natural language description of the initial configuration of a set of different colored blocks and determining how to leverage a small set of possible actions in order to achieve a goal configuration. Valmeekam et al. [57] directly compare models to non-expert humans, and demonstrate that humans can find a plan solution for 78% of the planning problems, and 89.7% of those plans were optimal (using the fewest number of steps). In contrast, the best-performing model (GPT-4) can find a plan solution for 35.6% of planning problems when leveraging chain-of-thought prompting and allowing the model to provide the plan in natural language. When the authors shift from the traditional description of Blocksworld problems to a ‘Mystery’ Blocksworld, wherein terms like “blue block” are consistently substituted for new, unrelated words or alphanumeric designations, zero-shot performance drops to 0% and one-shot performance to 2%. This research (along with later investigations [59, 58]) demonstrates that while the traditional Blocksworld problems are likely well-described on the web and therefore in LLM pretraining data, the Mystery condition reveals that successful performance does not reflect any actual mastery of the planning concepts that can be applied to a new set of items, even though the underlying structure of the problem states and actions hasn’t changed at all.
Furthermore, it has been shown that model performance drops consistently when their evaluation is shifted from the ‘default’ variant of a task to a ‘counterfactual’ (plausible, but not the default) variant. This provides even more compelling evidence that LLM ability to generalize information from a familiar task to a novel setting is quite limited. Wu et al. [67] test the largest models currently available across a battery of counterfactual tests, including arithmetic tests that shift to base-9 operations and natural language inference, in which models are prompted to indicate if premises that are both consistent and inconsistent with commonsense entail a series of conclusions. The authors show a clear and convincing drop in performance of models across counterfactual variants. Although the models perform better than chance, the authors conclude that LLM abilities are supported by non-transferable, default-condition-specific behaviors, rather than generalizable reasoning skills. We argue that this illustrates the limitation of models to Understanding capabilities, which are more superficial and rely on exemplification and classification, as opposed to Apply capabilities, which are precisely what is captured by these counterfactual tests in determining that models cannot generalize knowledge to a novel task setting. Importantly, it has been demonstrated that these shortcomings persist in the recent ‘reasoning’ models including GPT-o1 and DeepSeek-R1.
We emphasize that the drop in performance in these cases can be easily explained through our perspective, which posits that LLMs rely on context-guided extrapolation from pre-training priors. Specifically, counterfactuals, along with the types of questions posed in the faux pas test, are exactly the kind of scenarios wherein there are likely to have fewer relevant priors in the pre-training data.
4.2 Hallucinations
The second major piece of evidence against the notion of advanced reasoning in LLMs is their tendency to generate output that is not aligned with reality, a phenomenon known as hallucination [26]. This has also been referred to as confabulation and compared to human confabulation [55]. However, it is important to distinguish between the two, as there is no evidence to suggest that LLMs possess any agency (further discussed in Sections 5.1, 5.2.3 and 7), and because hallucinations can always be traced to the generation of a statistically most likely next token or pattern [24]. In fact, this second factor reinforces our perspective that LLMs are relying on priors from pre-training data guided by the context presented in the prompt. When the prompt lacks the necessary context to direct the model to the correct priors, LLM outputs default to the broader pre-training data (which determine the statistically most likely output), resulting in hallucinations.
5 Our View: Contextual Extrapolation from Training Priors
LLMs are unique in their ability to solve a broad range of tasks, even when trained primarily on next-token prediction [53]. Following the initial pre-training process, LLMs are further trained to ensure that they: a) follow instructions provided in their prompts, via instruction fine-tuning [61], and b) align their output with human preferences to mitigate the generation of harmful and offensive content [42]. We will first focus our discussion on the notion of ‘emergent’ advanced reasoning capabilities in LLMs that are not further trained beyond the initial pre-training (Base models), before then addressing the implications of post-training itself (instruction-tuned models).
5.1 Base Models: What Does Pre-Training Enable?
In exploring the complex reasoning capabilities of LLMs that are not post-trained (i.e. Base LLMs), it is well established that Base LLMs can generate responses beyond the next most likely token only when prompted with in-context examples [36]. Therefore, it is crucial to examine the nature of ICL and whether ICL, by itself, can lead to the emergence of advanced reasoning capabilities. Recall that ICL is the ability of Base LLMs to complete a task based on a few examples included in the prompt as exemplified by the modified arithmetic example presented in Section 1. In this section, we explore the question of whether ICL could, in a sufficiently large model, lead to advanced reasoning.
Olsson et al. [40] were the first to systematically explore the capabilities of LLMs in performing ICL with random tokens. They demonstrated that models can consistently complete patterns presented in their prompts, even when the patterns involve random tokens. Specifically, they showed that when presented with a sequence [A][B][A][A][B]...[A], LLMs could consistently respond with $[B]$, even if $A$, $B$, and the intermediate tokens were all random. The use of random tokens in this experiment was intentional and significant. The fact that the sequences were entirely random suggests that the models were not relying on memorization. This provides compelling evidence against the notion that LLMs are merely stochastic parrots.
This work, which has led to research exploring circuit discovery in LLMs [17], may give the impression that not only are LLMs not stochastic parrots, but they also exhibit some algorithmic capabilities. However, a key piece of recent evidence challenges this notion [39]. Recent work has shown that, while LLMs can complete sequences as described earlier, their ability to do so diminishes when the random tokens chosen are less frequent in pre-training data. This insight is crucial, as it highlights that these purported algorithmic capabilities are still closely tied to the pre-training data.
An important aspect of interpretability research is the observation that this tendency to repeat patterns can be seen at intermediate levels of LLMs [40]. This demonstrates that the previously mentioned $A$ and $B$ need not represent a single token but can instead correspond to entire sequences (such as a movie review) or concepts (such as sentiment [39]). We contend that viewing the problem through this lens enables us to understand how LLMs might extend their pattern completion capability to more complex tasks, such as labeling movie reviews with the appropriate sentiment. The idea that the ability to repeat patterns generalizes to the task-solving capabilities presented by ICL is further supported by work showing that ICL works even when labels are flipped [64]. Consider the example where the input consists of movie reviews, and the output is the sentiment of the review. In the standard ICL setting, we present examples of multiple movie reviews and their corresponding sentiments, and then provide an unseen movie review for the LLM to respond with the appropriate sentiment. Flipped-label ICL involves switching the positive and negative labels. Interestingly, ICL continues to be effective even after flipping the labels. Indeed, it has been demonstrated that replacing sentiment labels with semantically unrelated labels, such as ‘Foo’ and ‘Bar,’ still does not affect the ability of models to classify the movie reviews according to the new labels. This suggests that the pattern-completion mechanism observed earlier in the template-based completion remains central to how models solve tasks using ICL.
Does this imply that LLMs can use ICL to perform complex reasoning? We argue that this is not the case for two reasons: first, as mentioned earlier, ICL is still limited by the data used for pre-training, and we explore research indicative of this observation below. More importantly, however, the nature of ICL, which requires the inclusion of examples, means that models will always be constrained to performing tasks in ways that closely resemble the in-context examples provided [36]. This inherent characteristic of ICL suggests that LLMs cannot use it to perform advanced reasoning in ways that are not already clearly demonstrated by a user. Such lack of generalization fundamentally fails the requirements for advanced reasoning akin to human cognition (i.e. the human is doing the Apply process by pointing out the relevant generalization over patterns to be made based on the examples provided; the model need only Understand at a basic level the generalization made in the examples). Importantly, LLMs will never gain ‘agency’ in the ICL setting as what they do is always limited to user input through in-context examples.
Extrapolation from pre-training priors We derive the notion that ICL can be viewed as solving the task by use of priors extrapolated from pre-training data from the several theories put forward on how ICL functions. Indeed, a line of research directly suggests that ICL is influenced by the distributions within the pre-training data  [14, 23]. Similarly, work by Li et al. [34] proposes a framework based on Probably Approximately Correct (PAC) learning to explain ICL, while Zhang et al. [73] and  Xie et al. [68] suggest that ICL can be explained in terms of Bayesian inference. We note that both these explanations also imply a reliance on pre-training priors.
Additional research into ICL capabilities includes the work of Dai et al. [18], who suggest that ICL in LLMs may share similarities with fine-tuning, which is known to be a method of transfer learning enabling models to leverage pre-training priors to solve the downstream task they are being fine-tuned on. Other studies have shown that ICL might implicitly perform gradient descent and construct a function at inference time, particularly for regression tasks [1, 33, 72]. This has been linked to gradient-based meta-learning, as proposed by Von Oswald et al. [60]. Both of these explanations would, as with fine-tuning, imply a reliance on pre-training priors.
Importantly, we note that regardless of the specific underlying mechanism that makes ICL possible, all existing research points to a central thesis: ICL, whether it involves direct extrapolation from pre-training data, or has an equivalence to meta-learning, fine-tuning, or PAC learning, must rely directly on priors from pre-training data. Specifically, we can consider the in-context examples providing a way of guiding which of the pre-training priors are relevant to the task at hand, just as in the case of transfer learning through fine-tuning.
5.2 Post-Trained Models: What Does Post-Training Enable?
Thus far, we have discussed how models that are only pre-trained, without instruction tuning or other forms of post-training, can generate content that goes beyond merely predicting the statistically most likely next token, but only when ICL is applied. We have also explained why ICL itself does not constitute advanced reasoning of the type we are concerned with. Instead, we have shown that in these cases, the more likely explanation is that models are extrapolating based on priors extracted from pre-training, with those priors being guided by in-context examples. In this section, we will address the question of models that have undergone post-training, first through instruction tuning, and second through other post-training methods[61], including alignment to human preferences [42] and the more recent methods used in what are called ‘thinking models,’ which generate reasoning traces, such as Group Distributional Preference Optimization (GDPO) [70]. However, before discussing the inherent abilities of LLMs, it is important to first clarify what instruction tuning and similar processes mean in terms of task-specific training. This distinction complicates the evaluation of inherent abilities.
5.2.1 The Impact of Fine-Tuning
Unlike in the case of Base models, the evaluation of instruction-tuned models becomes more complex. These models are trained on a range of tasks, so determining whether they demonstrate emergent reasoning abilities requires evaluating whether the tasks they were trained on overlap with the tasks they are being tested on [7].
To clarify, we do not deny that LLMs’ ability to solve tasks increases as their scale grows, as argued by Kaplan et al. [28]. However, the ability of a model to understand, for example, social situations can be evaluated through tasks that involve questions about such scenarios. If the model performs above a random baseline on these tasks, we might describe this as an emergent ability. For instance, a model trained to answer questions about social situations at work that can then answer questions about social situations at home might exhibit generalization, but this alone would not qualify as ‘emergence’ of reasoning on the level that would imply advanced reasoning akin to human-like cognitive processes beyond the simplest levels of Remember and Understand. However, if a model is explicitly trained to answer questions about social situations, it may be able to apply this knowledge within a narrow domain, but generalizing to novel situations without clear out-of-domain performance would not constitute emergent reasoning. If the tasks are similar, the model’s performance is less indicative of emergent reasoning [7]. Understanding the degree to which instruction-tuned models can generalize to novel tasks is essential in assessing their true emergent abilities. Importantly, generalization over such a narrow domain is less indicative of the kind of high level cognitive capabilities that is of interest to us. Furthermore, recent research has demonstrated that, when the specific tasks used in instruction-tuning are taken into account, there is a statistically significant correlation between Base and instruction-tuned model performance [7] indicating that instruction tuning itself does not significantly alter the ability of LLMs.
5.2.2 Chain of Thought and ‘Thinking’ Models
Chain of Thought (CoT) prompting is a technique where LLMs are required to generate a rationale in addition to the answer, and it has been found to be highly effective [63]. Instruction tuning of models now routinely includes CoT data in their training [30], which is why most model outputs are verbose. However, recent research has shown that CoTs are not always faithful [27], and they can often generate incorrect reasoning traces while still reaching the correct answer[56]. Subsequent studies have suggested that even incorrect [54] reasoning traces can improve model performance, indicating that CoT is more about guiding the model through the solution space rather than producing fully rational reasoning. This notion is further reinforced by the work by  Prystawski et al. [45], who find that the effectiveness of the reasoning steps is based on the ‘local’ statistical distribution of pre-training data.
This suggests that methods built on CoT, such as GDPO [69], which loosely uses reinforcement learning to guide models towards better reasoning traces, such as GPT-o1 [74] and DeepSeek R1 [19] (also called ‘reasoning models’), still rely on approaches that are not fundamentally different from what we have discussed. Therefore, while these methods may provide more effective ways of guiding the model, they do not fundamentally differ from extrapolation from data priors directed through the context used within the prompts. Indeed, this notion is reinforced by the fact that models that ‘reasoning models’ continue to exhibit hallucinations [41] and continue to fail to generalize just as older models did [50].
5.2.3 Explaining Emergent Abilities through Contextual-Extrapolation
While there is a body of research showing that models of a sufficiently large size show unexpectedly sharp performance increases on benchmarks [21, 62], there is also significant debate surrounding emergent abilities, including criticism of the specific evaluation metrics used [48]. A detailed exploration of these purported ‘emergent abilities,’ by Lu et al. [36] offers several insights that are highly relevant to our discussion: first, they demonstrate that base models cannot solve benchmark tasks without the inclusion of in-context examples in the prompt. While instruction-tuned models, can solve a range of tasks, they argue that the ability of such post-trained models to respond to tasks may be driven by ICL, albeit in a form triggered without explicit examples made possible through instruction tuning [36]. This is demonstrated through a battery of 1,000 experiments showing that the tasks solvable by ICL and those solvable without examples by instruction-tuned models overlap significantly, indicating a common underlying mechanism. Using this framework, the authors explain emergent abilities as ICL becoming ‘more powerful’ (more expressive power) as models scale up, eventually reaching a critical point where they can solve specific tasks. This notion is further reinforced by recent work in mechanism interpretability by  Niu et al. [39], who use interim checkpoints of Pythia to demonstrate an increase in ICL ‘power’ with greater pre-training and larger model sizes. All of this evidence directly supports our view that context-directed extrapolation from pre-training priors is indeed what drives LLM performance, and can also explain the purported ‘emergent abilities’ in LLMs
6 Context Directed Extrapolation vs Stochastic Parroting
We would like to acknowledge that Context Directed Extrapolation and Stochastic Parroting share some commonalities. Therefore, this section clarifies the commonalities and differences between the stochastic parroting perspective and the perspective of context-directed extrapolation from training data priors.
Commonalities. From the perspective of base model performance, there is no functional difference between context-directed extrapolation and stochastic parroting. As previously discussed, base models consistently fail on tasks requiring reasoning when they are presented without examples. One can argue that the examples simply form a long context, making the correct answer the most probable sequence completion. This makes both theories appear to describe the same mechanism: the model completes a given sequence based on statistical patterns. Consequently, when analyzing base models alone, the two views are indistinguishable.
Differences. The functional divergence between these two views becomes apparent with instruction-tuned models. While a base model’s output can be seen as the most probable completion of a long sequence of examples, the context-directed extrapolation framework posits that instruction-tuning enables a different mechanism. Our framework suggests that instruction-tuning allows a model to interpret an instruction as a directive to construct an implicit task context, rather than as a literal sequence to be continued. This context activates relevant priors from the model’s pre-training data, enabling it to solve the specified task. The critical evidence supporting this distinction is that instruction-tuned models can solve logical deduction problems without any examples. This phenomenon is directly accounted for by context-directed extrapolation but cannot be explained by a strict stochastic parroting model. This functional gap becomes even more stark in tasks involving novel words, as described previously.
Grounding An important implication of context-directed extrapolation is that it facilitates a limited form of grounding. This is not grounding in the sense of connecting language to embodied experience, but rather a mechanism by which the model can access information that is not explicitly present in the surface text. Because the process involves extrapolating from priors activated by the prompt, a model can, for instance, take the definition of a nonce verb and productively apply that new meaning in novel contexts.
Therefore, while there is some overlap, our context-directed perspective is fundamentally distinct from a stochastic parroting mechanism, which by definition is incapable of handling novel inputs that fall outside its memorized distribution. The demonstrated ability of LLMs to extend prompt-based definitions, apply abstract patterns, and generate context-appropriate interpretations indicates that extrapolation provides access to information not reducible to surface statistics. In this sense, context-directed extrapolation offers a framework to explaining LLM capabilities and limitations, including a form of limited grounding, which remain contingent upon the model’s training priors and the context supplied at inference.
7 Conclusions & Recommendations
The notion that LLMs have unpredictable, emergent reasoning capabilities has motivated the possibility that LLMs possess either consciousness, agency, or both, which has, in turn, raised safety issues around preparing for this possibility [35]. While we commend both developer and policymaker efforts to prepare for a range of possible AI advances, we propose that near-term research be guided by the more realistic view that current LLM capabilities should be understood as context-directed extrapolation. This perspective will help alleviate fears that LLMs are ‘deceptive,’ as studies into the deceptive capabilities of LLMs simulate intent through prompts under the assumption that such agency will eventually become apparent in the future [22, 44, 49, 29].
Additionally, when viewed as context-driven extrapolation from training priors, including priors from post-training processes such as instruction tuning and GDPO, several previously perplexing aspects of LLMs become clear and explainable. For example, we can now understand why LLMs require carefully curated prompts and why they tend to fail without them. This is because the specific prompt used with instruction-tuned models will determine the models’ ability focus on the relevant priors.
Similarly, it becomes clear why models can struggle with counterfactuals, even simple ones that children can solve, such as the faux-pas test (Section 4.1), while still being able to tackle more complex problems of a similar nature. For example, while LLMs fail on social intelligence tasks, including Theory of Mind in the faux-pas test, they succeed on the more traditional and well-known “Sally-Anne” test [4] (in which Sally places an object in one location, Anne moves it while Sally is out of the room, and then asks where Sally will expect the object to be). This results in conflicting outcomes on Theory of Mind evaluations. Therefore, we recommend using counterfactuals to assess true inherent abilities, while employing more standard test sets to evaluate cases where the model’s ability to solve a task is of primary interest, regardless of whether it relies on priors.
Finally, understanding the role and importance of context-directed extrapolation in LLMs has the potential to facilitate more effective post-training methodologies that will advance capabilities in a manner that is fully controllable and predictable, increasing trust in LLMs.
In this work, we aim to provide a framework for understanding how LLMs function, grounded in existing literature. We demonstrate that our framework, which posits that LLMs respond through context-driven extrapolation from training data priors, adequately explains both the capabilities and limitations of LLMs. We position our work within the context of two dominant narratives about these models: one that asserts LLMs are merely ‘stochastic parrots,’ and another that suggests they are either approaching or already exhibit advanced reasoning capabilities akin to high-level human cognition. While our work addresses a range of elements associated with LLMs, it does not cover every aspect, including future methods and, for example, those involving agentic AI. We contend that any evidence of advanced reasoning in these or future scenarios must account for context-driven priors. Additionally, we do not claim that all future AI poses no threat. AI safety preparations should be made, but the research outlined here demonstrates that such preparations should be grounded in understanding the clearest current source of LLM capabilities: context-directed extrapolation.